# I know What You Sync: Covert and Side Channel Attacks on File Systems via syncfs

### v1

@@ 1 Introduction
Security and privacy issues in multi-user operating systems (OS) and application software have drawn significant attention from both academia and industry over the past decades. These systems enable concurrent access by multiple users, facilitating efficient sharing of computing resources at both hardware and software levels. At the hardware level, research has identified a series of side-channel attacks targeting shared CPU caches [1, 2, 3], hardware sensors (e.g., accelerometers, camera and gyroscope) [4, 5, 6], DRAM [7, 8], interconnects [9, 10], GPUs [11, 12], FPGAs [13, 14] and more. On the software side, numerous studies have revealed security vulnerabilities in public OS-provided APIs, leading to OS-level side-channel attacks on systems like iOS [15, 16], Android [17, 18], and Linux [19, 20].
@@ 2 Background and Threat Model
This section begins with an overview of the Linux file system and I/O buffers. Next, we explain the key system calls relevant to I/O operations. Finally, we present our threat model.
@@ 2.1 Linux File System
The next layer consists of the actual file systems and their caching mechanisms. These include the inode cache, which stores file metadata, the page cache, which buffers file data, and the journal cache, which tracks transactional logs to ensure data integrity. These caches improve performance by minimizing direct disk access, allowing the system to process file operations more quickly.
Figure 1: Overviw of Linux file system and I/O system calls between the application layer and the physical device.
2.2 Linux I/O Buffers
Inode cache. The inode cache is a memory-resident structure in the main memory that stores recently accessed inodes, which are data structures containing metadata about files and directories. Its purpose is to bypass disk reads and writes, thereby reducing latency and speeding up common I/O operations. This caching mechanism plays a critical role in optimizing file system performance.
@@ 3.2 Inferring write Size via syncfs
Experiment 3: infer write size via syncfs. We have two steps for this experiment: within page size (4KB), and above page cache size. Within page size, we set the initial write size as 64 bytes. It keeps increasing with a stride of 64 bytes and ends at 4 KB. Above page size, we set the initial write size as 4 KB. It keeps increasing with a stride of 4 KB and ends at 64 KB. The experiment was conducted on an NVIDIA Jetson AGX Orin machine with the ext4 file system set to default and running kernel version 5.15.136-tegra.
@@ 5.1 Attack 1: Web Fingerprinting on Linux
Comparison with Sync+Sync [20]. Jiang et al. [20] introduced fsync-based contention leakage for website fingerprinting, requiring both the victim and attacker to use the fsync system call to create contention. Their approach, however, is limited by the similar number of fsync calls across websites, resulting in low accuracy for website fingerprinting, with reliable identification for only two websites (qq.com and sina.com.cn). In contrast, our method utilizes each website’s unique syncfs delay pattern, which reflects specific write frequencies and content sizes. It imposes no special requirements on the victim user’s behavior (e.g., fsync calls). Our approach achieves over 93% average F1 score for 100-class website fingerprinting.
@@ 5.2 Attack 2: Video Fingerprinting on Linux
In our attack, we assume the victim uses ffplay, a lightweight media player based on FFmpeg libraries, to watch video streams, while a separate host machine runs ffmpeg as the streaming server. Meanwhile, the attacker profiles the write patterns of the victim’s ffplay in the background. For smooth playback, ffplay uses internal buffering for network streams, with different video content producing distinct write patterns in this buffer. Consequently, the attacker can fingerprint the victim’s video by profiling these write patterns via the syncfs system call.
@@ 5.3 Attack 3: Application Fingerprinting on Android
TABLE VI: Applications evaluated for the fingerprinting attack.
@@ 7 Potential Mitigations
Detecting abnormal syncfs usage. Our attack relies on continuous probing of syncfs timing delays, and a potential defense could involve monitoring for abnormal syncfs query patterns, such as using a daemon process to track suspicious syncfs activity. Previous research [67, 68] has explored malicious application detection by monitoring API call sequences and return values.
@@ 8 Related Work
I/O-related attacks. Our attacks fall into the category of I/O-related attacks. Prior research has demonstrated that contention leakages on I/O can be exploited to establish covert and side channels. On one hand, timing leakages from contention have been shown on physical I/O components in modern computers, such as PCIe, CPU ring bus, and NVLink. For example, Tan et al. [9] introduced PCIe congestion side-channel attacks across GPUs, Network Interface Cards (NICs), and Solid-State Drives (SSDs). Similarly, Mert Side et al. [72] introduce a novel side channel vulnerability exploiting contention on the host-GPU PCIe bus, a critical interface in modern GPUs used in both traditional and cloud computing. Paccagnella et al. [73] propose the first microarchitectural side channel attacks leveraging contention on the CPU ring interconnect. Similar contention-based attacks have also been executed on other hardware architectures, such as CPU Mesh [74]. Recently, Zhang et al. [10] reveal vulnerabilities in NVLink interconnects, demonstrating a covert channel across GPUs and an application fingerprinting side channel attack using congestion-based timing leakages.
On the other hand, software I/O has also been investigated for contention-based leakages. Gruss et al.[75] introduced a software side-channel attack on the OS page cache, allowing unprivileged monitoring of memory access patterns. More recently, Jiang et al.[20] presented Sync+Sync, a timing-based covert channel leveraging contention in fsync calls on shared persistent storage. Similarly, Chen et al. [19] proposed a set of software cache write covert channels that exploit file system synchronization primitives.
Both of these works depend on contention from concurrent fsync API invocations and are limited to forming covert channels. Since most applications do not use the fsync API, this significantly restricts their potential for effective side-channel attacks. In contrast, our work exploits delay patterns in syncfs, a fundamentally different approach to contention-based leakages. Our method does not require the victim application to execute any I/O system calls, enabling us to implement both robust covert channels and high-resolution side channel attacks.

### v2

@@ 1 Introduction
Security and privacy issues in multi-user operating systems (OS) and application software have drawn significant attention from both academia and industry over the past decades. These systems enable concurrent access by multiple users, facilitating efficient sharing of computing resources at both hardware and software levels. At the hardware level, research has identified a series of side-channel attacks targeting shared CPU caches [1, 2, 3], hardware sensors (e.g., accelerometers, camera and gyroscope) [4, 5, 6], DRAM [7, 8], interconnects [9, 10], GPUs [11, 12], and more. On the software side, numerous studies have revealed security vulnerabilities in public OS-provided APIs, leading to OS-level side-channel attacks on systems like iOS [13, 14], Android [15, 16], and Linux [17, 18].
@@ 2 Background and Threat Model
This section overviews the Linux file system as well as the key system calls relevant to I/O operations. Finally, it presents the threat model assumed in the paper.
@@ 2.1 Linux File System
The next layer consists of the actual file systems and their caching mechanisms. These include the inode cache, which stores file metadata, the page cache, which buffers file data, and the journal cache, which tracks transactional logs to ensure data integrity. These caches improve performance by minimizing device access, as we elaborate below.
Figure 1: Overview of Linux file system and I/O system calls between the application layer and the physical device.
Inode cache. The inode cache is a memory-resident structure in the main memory that stores recently accessed inodes, which are data structures containing metadata about files and directories. Its purpose is to bypass device reads and writes, improving I/O performance. This caching mechanism plays a critical role in optimizing file system performance.
@@ 3.2 Inferring write Size via syncfs
Experiment 3: infer write size via syncfs. We have two steps for this experiment: within page size (4KB), and above page cache size. Within the page size, we set the initial write size as 64 bytes. It keeps increasing with a stride of 64 bytes and ends at 4 KB. Above page size, we set the initial write size as 4 KB. It keeps increasing with a stride of 4 KB and ends at 64 KB. The experiment was conducted on an NVIDIA Jetson AGX Orin machine with the ext4 file system and running kernel version 5.15.136-tegra.
@@
3.3 Robustness of Side Channel against Noise
In real-world settings, background applications may run concurrently with an attacker, introducing noise that can interfere with the side-channel. To assess the impact of such interference, we use the signal-to-noise ratio (SNR), a standard metric to evaluate side-channel leakage [37]. Specifically, we apply the following SNR equation: $SNR=\frac{Var(Signal)}{Var(Noise)}$. In particular, an SNR value greater than 1 is exploitable as recent work highlighted [16], with higher values indicating a higher quality channel. We collect the syncfs signals of browsing a website while introducing background noise in two ways: (1) opening additional browser tabs and (2) running I/O noise microbenchmarks. The I/O noise microbenchmarks launch multiple threads to continuously generate dirty pages at a rate of 81 MB/s. We computed the SNR by averaging the results over 10 traces. We observe that the SNR drops from 15.77 (without noise) to 15.54 and 12.78 when one and five browser tabs are opened, respectively. In a different experiment, when one and five I/O noise workloads are introduced to run concurrently with our spy and victim, the SNR decreases to 7.66 and 4.75, respectively.
@@ 5.2 Attack 2: Video Fingerprinting on Linux
In our attack, we assume the victim uses ffplay, a lightweight media player based on FFmpeg libraries, to watch video streams, while a separate host machine runs ffmpeg as the streaming server. Meanwhile, the attacker profiles the write patterns of the victim’s ffplay in the background. For smooth playback, ffplay uses a file system-backed temporary file to store received raw data. As the MPEG video is encoded, the received data is written to this file. The size and frequency of these writes are closely correlated with the visual content and dynamics of the video. Higher-resolution videos yield higher-quality signals, as the encoder operates more slowly and produces larger outputs, resulting in richer side-channel traces and improved classification accuracy. Consequently, the attacker can fingerprint the victim’s video by profiling these write patterns via the syncfs system call.
@@
Web vs video. We assume, consistent with prior web and video fingerprinting attacks, that the attacker already knows whether the victim is browsing the web or watching videos. However, this assumption is not critical since we observe that syncfs exhibits distinct timing patterns across these activities that would allow an attacker to recognize when these applications execute. Thus, a spy could continuously collect syncfs traces, which are then processed to identify when our target applications start to spy on them. To provide evidence that the signal can be used to identify applications, we show that we can tell traces from our two target applications apart. In particular, we train a binary classifier using the same CNN model employed in our fingerprinting attacks to infer the victim’s activity type, to provide evidence that the attack does not rely on knowing when the target applications are launched. The dataset contains 4,000 signal traces, evenly split between web browsing and video watching. Our model achieves an F1 score exceeding 99% across 10-fold cross-validation, with a standard deviation of just 0.12%. The average precision and recall are 99.98% and 99.82%, respectively, with standard deviations of 0.07% and 0.25%.
Robustness against noise. Our preceding experiments were conducted in a relatively quiet environment. However, in real-world settings, other applications may also run on the same file system, introducing noise that degrades the quality of side-channel leakages. To evaluate the impact of such noise on the video fingerprinting attack, we ran a web browser in the background and opened additional browser tabs, then re-evaluated the attack under these conditions. We used the same experimental testbed and classifiers for consistency. As a result, when one browser tab was opened in the background, the average F1 score on YouTube dropped from 95.69% to 79.57%, with a standard deviation of 12.61%. When five tabs were opened, the F1 score dropped further to 74.93%, with a standard deviation of 13.75%. Even with noise, the results remain far above the 5% random-guess baseline for 20-class classification. This demonstrates that the side channel remains leaky and exploitable even in noisier environments.
@@ 5.3 Attack 3: Application Fingerprinting on Android
TABLE VI: Applications evaluated for the application fingerprinting attack.
@@ 7 Potential Mitigations
Detecting abnormal syncfs usage. Our attack relies on continuous probing of syncfs timing delays, which makes it a promising target for detection. A potential defense is to monitor abnormal syncfs invocation patterns, such as periodic or high-frequency calls, by using a daemon process that tracks syncfs activity. Prior research [66, 67] has demonstrated the effectiveness of detecting malicious applications through API call sequence and return value monitoring, and similar techniques [68] could be applied here. Detection-based approaches may result in harmful false positives for applications that may naturally require higher frequencies of syncfs. Moreover, to evade such detection, an attacker may attempt to reduce the syncfs query rate, trading stealth for reduced signal quality. As shown in Figure 10, we reduce the sampling rate from the maximum achievable values (150 kHz on Android and 120 kHz on Linux); this reduction comes at the cost of decreased attack success rate, but an attacker may still glean useful information while avoiding detection.
@@ 8 Related Work
I/O-related attacks. Our attacks fall into the category of I/O-related attacks. Prior research has demonstrated that contention leakages on I/O can be exploited to establish covert and side channels. On one hand, timing leakages from contention have been shown on physical I/O components in modern computers, such as PCIe, CPU ring bus, and NVLink. For example, Tan et al. [9] introduced PCIe congestion side-channel attacks across GPUs, Network Interface Cards (NICs), and Solid-State Drives (SSDs). Similarly, Mert Side et al. [72] introduce a novel side channel vulnerability exploiting contention on the host-GPU PCIe bus, a critical interface in modern GPUs used in both traditional and cloud computing. Paccagnella et al. [73] and Dutta et al. [74] propose microarchitectural side channel attacks leveraging contention on the CPU ring interconnect. Similar contention-based attacks have also been executed on other hardware architectures, such as CPU Mesh [75]. Recently, Zhang et al. [10] reveal vulnerabilities in NVLink interconnects, demonstrating a covert channel across GPUs and an application fingerprinting side channel attack using congestion-based timing leakages. On the other hand, software I/O has also been investigated for contention-based leakages. Gruss et al. [76] introduced a software side-channel attack on the OS page cache, allowing unprivileged monitoring of memory access patterns.
Comparison with fsync-based attacks. More recently, Chen et al. [17] proposed a set of software cache write covert channels that exploit file system synchronization primitives. Similarly, Jiang et al. [18] presented Sync+Sync, a timing-based covert channel leveraging contention in fsync calls on shared persistent storage. Both of these works depend on contention from concurrent fsync API invocations and are limited to forming covert channels. Jiang et al. [18] introduced fsync-based contention leakage for website fingerprinting, requiring both the victim and attacker to use the fsync system call to create contention. However, their approach is limited by the similar number of fsync calls across websites, resulting in low accuracy for website fingerprinting, with reliable identification for only two websites (qq.com and sina.com.cn). In contrast, our method utilizes each website’s unique syncfs delay pattern, which reflects the sequence of write operations and sizes, patterns that are unique to the website content and structure. As a result, the attack achieves over 93% average F1 score in a 100-class website fingerprinting attack.
@@ 9 Concluding Remarks
Acknowledgements
This research was sponsored in part by the OUSD(R&E)/RT&L and was accomplished under Cooperative Agreement Number W911NF-20-2-0267. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARL and OUSD(R&E)/RT&L or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. We would like to thank the paper reviewer and, especially, the shepherd whose comments and requests substantially improved the paper.
Appendix A Meta-Review
A.1 Summary
This paper introduces a new class of side-channel attacks exploiting shared file system code. It identifies the syncfs system call as a new source of leakage that enables an attacker to infer activities of other processes on the system.
A.2 Scientific Contributions
  • Identifies an Impactful Vulnerability
  • Provides a Valuable Step Forward in an Established Field
A.3 Reasons for Acceptance
  1. 1.
    This paper identifies an impactful vulnerability in file system code that is present in virtually all existing Linux systems. It demonstrates how to carefully control and leverage syncfs delays to instantiate covert channels and infer behaviors of victim processes.
  2. 2.
    The paper provides a valuable step forward in an established field. Secure file and storage systems have been well-studied, but recent security research has shown that side-channel attacks are present at virtually all layers of the software stack. This paper highlights that even core system software is vulnerable to side-channel attacks and provides some insights on potential mitigations.
A.4 Noteworthy Concerns
  1. 1.
    The attacks rely on high sampling rates (150K calls per second). Attacks therefore have anomalous behavior that might be easily detected by simple (or state of the art) anomaly detection techniques. The takeaways could be stronger if the paper evaluated the detectability of the attack to some extent.
Appendix B Response to the Meta-Review
We thank the reviewers and the shepherd for their insightful comments. Current defenses do not detect the attack. We acknowledge that anomaly detection mechanisms might be configured to flag high syncfs() rates. However, it is not clear whether a low detection threshold is feasible without introducing false positives or whether the attacker can adjust the sampling rate to remain below the detection threshold while still exploiting the side channel.